# SimuRA: Towards General Goal-Oriented Agent via Simulative Reasoning Architecture with LLM-Based World Model

### v1

@@ 3.4 Design of Simulative Reasoning Agent Using LLM-Based World Model
The dominant approach to encoding observation $o_{t}$ (e.g., webpages, video streams) has been to directly pass all input tokens into an LLM to form continuous embeddings $\hat{s}^{z}_{t}$. While technically preserving all information, real-world sensory readings often suffer from inherent noise and high variability (e.g., ads on a webpage, varying weather and lighting conditions in video), which can make them brittle for reasoning over. Human cognition, on the other hand, has evolved to counter this variability by categorizing raw perception into discrete concepts barrett2017emotions , which are often encoded in language, symbols or structured thoughts. Indeed, natural language is inherently hierarchical, capable of encoding concepts from concrete ones (e.g., apple) to highly abstract ones (e.g., religion). Discrete representations are also complete in general xing2025critiques , which ensures no information is necessarily lost in the compression process. Implementing this form of perception, we propose to represent the world state $\hat{s}_{t}$ using a discrete natural language summary $\hat{s}^{c}_{t}$ generated by an encoder LLM $h$, formally expressed as below:
@@
Where each $\hat{s}_{t,i}$ is a natural language token. Likewise, we also implement the WM $f$ using an LLM which predicts the next state $\hat{s}_{t+1}$ as a natural language sequence $\hat{s}^{c}_{t+1}$, formally as below:
@@
The customary approach to decision-making with world models has been to perform simulations or rollouts based on the specific action space $\mathcal{A}(\pi)$ afforded to the agent. While this approach indeed captures all the execution details, the specific idiosyncracies of individual action spaces (e.g., parameter ordering, format, and scale) may distract from transferring knowledge across different action spaces, environments, and tasks for generalizable reasoning. The real world may contain a richer range of intentions than what a particular action space offers (e.g., clicking on a flight may mean either exploring the pricing or committing to the option). Last but not least, the sequential roll-out over atomic actions can be inefficient and increase opportunities for error accumulation across multi-step, low-level predictions (e.g., swooshing of liquids with each muscle twitch), when higher-level dynamics over more abstract actions (e.g., spilling water due to tilting the glass) remain stable and predictable. To close this gap, we adopt a hierarchical architecture which separates high-level, flexible planning from low-level, rigorous execution sutton1991dyna . As illustrated in Figure 3, the agent’s policy $p_{\tilde{\pi}}(a^{\prime}_{t}\mid\hat{s}_{t})$ and world model $p_{f}(\hat{s}_{t+1}\mid\hat{s}_{t},a^{\prime}_{t})$ operate over simulated actions $a^{\prime}_{t}$ from a separate action space $\mathcal{A}^{\prime}$, while another actor $p_{\nu}(a_{t}\mid a^{\prime}_{t},\hat{s}_{t})$ is responsible for selecting the concrete action $a_{t}\in\mathcal{A}$ conditioned on the selected simulated action $a_{t}^{\prime}$. This divide-and-conquer approach allows for more generalized reasoning disentangled from the exact details of the concrete action space and enables representation of a richer set of intentions. Furthermore, each simulated action $a_{t}^{\prime}$ may represent multiple execution steps in the environment (e.g., “explore the website” vs “click on the link”), which shortens the number of rollout steps for higher efficiency and fewer chances for error accumulation. In practice, we represent simulated actions $a_{t}^{\prime}$ using natural language due to its generality and expressivity, and find it results in more diverse and grounded action proposals, leading to better task success.
@@
Having discussed our major designs, we proceed to describe the full decision process of the SimuRA architecture: As illustrated in Figure 3, given observation $o_{t}$ (e.g., webpage screenshots and/or accessibility tree), SimuRA first infers the world state $\hat{s}_{t}$ using the encoder $h$, and then selects the best simulated action $a_{t}^{\prime*}$ through the planner. Inside the planner, the architecture performs simulative reasoning by proposing actions $a_{t}^{\prime}$ using policy $\tilde{\pi}$ and predicting the next state $\hat{s}_{t+1}$ using the world model $f$, and evaluating goal progress $\sum_{k=t}^{T^{\prime}-1}\gamma_{k}r(g,\hat{s}_{k})+\gamma_{T^{\prime}}V_{\pi,f}^{g}(\hat{s}_{T^{\prime}})$ using critic $v$ upon reaching state $\hat{s}_{T^{\prime}}$ at the planning horizon $T^{\prime}$. This can repeat multiple times until the planner selects the action sequence $a^{\prime*}_{t:T^{\prime}-1}$ with the highest expected success and passes the first step $a_{t}^{*}$ to actor $\nu$ which finally outputs the concrete action $a_{t}$. Formally, SimuRA can be seen as solving the following multi-level optimization problem:
@@
$a_{t:T^{\prime}-1}^{\prime*}$ $={\underbrace{\operatorname*{arg\,max}_{a_{t:T-1}^{\prime}}}_{\scriptsize\shortstack{sampled from policy $\tilde{\pi}$}}}\ \sum_{\hat{s}_{t+1:T^{\prime}}}{\underbrace{v(\hat{s}_{T^{\prime}})}_{\text{critic}}}\prod_{k=t}^{T^{\prime}-1}\ {\underbrace{p_{f}(\hat{s}_{k+1}\mid\hat{s}_{k},a_{k}^{\prime})}_{{\scriptsize\shortstack{world model}}}}$ (Planning) (8)
$a_{t}$ $=\operatorname*{arg\,max}_{a}\ {\underbrace{p_{\nu}(a\mid\hat{s}_{t},a_{t}^{\prime*})}_{\begin{array}[]{c}\scriptsize\text{actor}\text{}\end{array}}}$ (Acting) (10)
In practice, we implement each of these components by zero-shot prompting pretrained LLMs. While these LLMs alone are often insufficient for many complex agentic tasks, SimuRA’s divide-and-conquer approach combines existing LLM strengths like instruction-following, summarization, reflection, and tool use to allow agentic behavior to emerge. Benefiting from massive web-scale pretraining on next-token prediction $p(x_{t}\mid x_{<t})$, which is formally akin to world modeling, LLMs possess significant potential to serve as world models with natural-language state and action spaces hao2023reasoninglanguagemodelplanning ; hu2023language . We approximately infer the world state $\hat{s}_{t}$ and action $a_{t}$ by sampling from the LLM-based encoder and actor distributions $p_{h}$ and $p_{\nu}$, respectively. For planning, we optimize over the sampled actions $a_{t:T^{\prime}-1}^{\prime}$ using readily available tree search algorithms like Depth-First Search (DFS) and Monte-Carlo Tree Search (MCTS).
Figure 4: Optimal agent architecture design with conditional probability annotation
@@ 4 Experiments
    BrowsingAgent from OpenHands  (openhands, ), a representative open-web agent which generates chain-of-thought before selecting an action
@@
Figure 5: LLM-based implementation of our proposed agent model for web-related tasks (e.g. multi-website QA, flight search, etc). Planner is where we implement our proposed world-model-based planning. We also implement a baseline that simply samples the plan from a language model (i.e., autoregressive planning).
Figure 5 presents our implementation when applied to web browsing. We use prompts tailored to the web environments in this example, but plan to extend to other environments and move towards training a single agent model that can act optimally in all environments, which is an exciting next step. At each step $t$, the agent receives the observation $o_{t}$ as the HTML-based accessibility tree visible through the browser’s viewport (an example is provided in Appendix A). The agent then uses encoder LLM $h$ to summarizes the observation as $\tilde{s}_{t}\sim p_{h}(\cdot\mid o_{t})$, and then add it to a selective memory of past summaries and simulated actions $\{m(\tilde{s}_{k},a_{k}^{\prime*})\}_{k=1}^{t-1}$ to form the estimated world state $\hat{s}_{t}=[m(\tilde{s}_{1},a_{1}^{\prime*}),\dots,m(\tilde{s}_{t-1},a_{t-1}^{\prime*}),\tilde{s}_{t}]$ for planning. During planning, we sample $M$ simulated actions $a^{\prime}_{t}$ from the policy $\tilde{\pi}$, cluster them into distinct actions, and use the world model $f$ to predict the next summary as $\tilde{s}_{t+1}\sim p_{f}(\cdot\mid\hat{s}_{t},a_{t}^{\prime})$ to form the next state $\hat{s}_{t+1}=[m(\tilde{s}_{1},a_{1}^{\prime*}),\dots,m(\tilde{s}_{t},a_{t}^{\prime}),\tilde{s}_{t+1}]$; this repeats until the planning horizon $T$. To evaluate the terminal state $\hat{s}_{T}$ with critic $v$, we prompt the LLM to generate qualitative answers and convert them into numerical scores (e.g., “success” receives a score of 1), and repeat for $N$ times to capture the fine-grained differences between states. Following previous work koh2024tree ; gu2025llmsecretlyworldmodel , we set $M=N=20$ and $T=t+1$, and use DFS as the search algorithm. We implement the planning process using LLM Reasoners hao2024llm , a library for LLM-based complex reasoning using advanced algorithms. After the planner selects the simulated action $a_{t}^{\prime*}$, we update the memory with $m(\tilde{s}_{t},a_{t}^{\prime*})$. For the actor $\nu$, we additionally include the observation text $o_{t}$ in the prompt to ensure the action grounding. All the prompts are included in Appendix B.
@@
An overview of our results is presented in Figure 6. Across all 3 categories of tasks, our architecture shows a clear advantage over the baseline BrowsingAgent, specifically increase the success rate on complex website navigation from 0% to 32.2%. Our proposed world model reasoning for planning also consistently improves over simple planning with autoregressive LLMs by up to 124%. In the subsections below, we will introduce the evaluation settings and discuss the results in more detail.
Figure 6: Overview of performance comparison between SimuRA and baselines. The full architecture shows clear advantage over the baseline BrowsingAgent, improving the performance on complex website navigation from 0% to 32.2%. Our proposed world model reasoning for planning also consistently improves over simple planning with autoregressive LLM by up to 124%.
@@ 4.1 Complex Website Navigation
Dataset
Due to the lack of benchmarks for complex website navigation on the open internet, we created our own dataset using flight search as an example, since it is an important part of travel planning, a major use case for web agents. We formulate flight search as the user asking for a flight satisfying a list of constraints (e.g., one-way, from New York to Los Angeles), and the agent operating a flight search engine to find such a flight and returning it to the user. To evaluate the agent’s reasoning ability, we further produce questions with varying number of constraints by iteratively adding to the list, which enables a counterfactual analysis that controls for the confounding effect of specific constraint configurations (e.g., an agent with perfect reasoning should still be able to answer the same question with one more constraint; an agent of rote memorization will likely fail when the question changes slightly).
Figure 8: Illustration of the data generation process for the FlightQA dataset. We first prompt a LLM to generate $N$ lists of $C$ starting constraints (Constraint Generation). Then, we prompt the LLM to iteratively add constraints to the lists one by one, repeating for $K$ times (Constraint Extension). Finally, we prompt the LLM to convert each constraint list into a question in natural language (Question Generation).
@@
We ran the experiments and evaluation using gpt-4o between November 24th, 2024 and December 9th, 2024. For the environment, we use BrowserGym (workarena2024, ), a popular open-source browser sandbox. We stop each run when the agent provides a response or after the agent takes 30 actions, whichever comes first. We also mark the run as failed when the agent repeats the same action for 3 times consecutively or when the agent causes more than 3 errors while interacting with the browser.
@@ Results
We present our Complex Website Navigation results in Table 1. Compared to BrowsingAgent which fails completely in this task, our full architecture improves the correct rate from 0% to 32.2%. Within our architecture, our proposed world-model-based planning shows superior performance over autoregressive reasoning with a 124% improvement (significant at the 0.01 level). The other components in our architecture, which communicate using the concept-based latent space of model-generated language (e.g., observation summary and selective memory), also result in more coherent behavior by reducing the action error rate in BrowsingAgent from 93.3% to 1.1%. However, the autoregressive reasoning still results in frequent repetitions, which is mitigated by world model-based planning (44.4% $\rightarrow$ 18.9%).
@@
Table 1: Performance and outcome statistics for the FlightQA dataset. Our architecture increases the correct rate from 0% in OpenHands BrowsingAgent to 32.2%. Reasoning by world model simulation also clearly outperforms autoregressive reasoning by 124%. ** indicates being significantly higher than the second-best method at the statistical significance level of 0.01 ($p<0.01$) based on pairwise t-test. We implement the autoregressive planner with o1 and o3-mini, respectively.
@@
We present our results on General Web Automation in Table 3. Continuing the patterns from previous experiments, our proposed architecture improves over BrowsingAgent by up to 91.7%, while within our architecture, world model reasoning improves over autoregressive reasoning by 21.1%, highlighting the comparative advantage under the given experimental setup.

### v2

@@ 3.4 Design of Simulative Reasoning Agent with Concept-based Latent States and Hierarchical Planning
The dominant approach to encoding observation $o_{t}$ (e.g., webpages, video streams) has been to directly mapping all input tokens into continuous embeddings with fixed dimensionalities $\hat{s}^{z}_{t}$. While technically preserving all information, real-world sensory readings often suffer from inherent noise and high variability (e.g., ads on a webpage, varying weather and lighting conditions in video), which can make them brittle to reason over. Human cognition, on the other hand, has evolved to counter this variability by categorizing raw perception into discrete concepts barrett2017emotions , which are often encoded in language, symbols or structured thoughts. Indeed, natural language is inherently hierarchical, capable of encoding concepts from concrete ones (e.g., apple) to highly abstract ones (e.g., religion). Discrete representations are also complete in general xing2025critiques , which ensures no information is necessarily lost in the compression process. Implementing this form of perception, we propose to represent the world state $\hat{s}_{t}$ using a discrete natural language summary $\hat{s}^{c}_{t}$ generated by a pretrained encoder model $h$, formally expressed as below:
@@
Where each $\hat{s}_{t,i}$ is a natural language token. Likewise, we also denote the WM $f$ which predicts the next state $\hat{s}_{t+1}$ as a natural language sequence $\hat{s}^{c}_{t+1}$, formally as below:
@@
The customary approach to decision-making with world models has been to perform simulations or rollouts based on the specific action space $\mathcal{A}(\pi)$ afforded to the agent. While this approach indeed captures all the execution details, the specific idiosyncrasies of individual action spaces (e.g., parameter ordering, format, and scale) may hinder the transfer of knowledge across different action spaces, environments, and tasks, thereby limiting generalizable reasoning. Indeed, the real world may contain a richer range of intentions than what a particular action space offers (e.g., clicking on a flight may mean either exploring the pricing or committing to the option). Last but not least, the sequential roll-out over atomic actions can be inefficient and increase opportunities for error accumulation across multi-step, low-level predictions (e.g., swooshing of liquids with each muscle twitch), when higher-level dynamics over more abstract actions (e.g., spilling water due to tilting the glass) remain stable and predictable. To close this gap, we adopt a hierarchical architecture which separates high-level, flexible planning from low-level, rigorous execution sutton1991dyna . As illustrated in Figure 3, the agent’s policy $p_{\tilde{\pi}}(a^{\prime}_{t}\mid\hat{s}_{t})$ and world model $p_{f}(\hat{s}_{t+1}\mid\hat{s}_{t},a^{\prime}_{t})$ operate over simulated actions $a^{\prime}_{t}$ from a separate action space $\mathcal{A}^{\prime}$, while another actor $p_{\alpha}(a_{t}\mid a^{\prime}_{t},\hat{s}_{t})$ is responsible for selecting the concrete action $a_{t}\in\mathcal{A}$ conditioned on the selected simulated action $a_{t}^{\prime}$. This divide-and-conquer approach allows for more generalized reasoning disentangled from the exact details of the concrete action space and enables representation of a richer set of intentions. Furthermore, each simulated action $a_{t}^{\prime}$ may represent multiple execution steps in the environment (e.g., “explore the website” vs “click on the link”), which shortens the number of rollout steps for higher efficiency and fewer chances for error accumulation. In practice, we represent simulated actions $a_{t}^{\prime}$ using natural language due to its generality and expressivity, and find it results in more diverse and grounded action proposals, leading to better task success.
Figure 4: Illustration of the design of simulative reasoning agent with concept-based latent states and hierarchical planning.
@@
Having discussed our major designs, we proceed to describe the full decision process of the SimuRA architecture: As illustrated in Figure 4, given observation $o_{t}$ (e.g., webpage screenshots and/or accessibility tree), SimuRA first infers the world state $\hat{s}_{t}$ using the encoder $h$, and then selects the best simulated action $a_{t}^{\prime*}$ through the planner. Inside the planner, the architecture performs simulative reasoning by proposing actions $a_{t}^{\prime}$ using policy $\tilde{\pi}$ and predicting the next state $\hat{s}_{t+1}$ using the world model $f$, and evaluating goal progress $\sum_{k=t}^{T^{\prime}-1}\gamma_{k}r(g,\hat{s}_{k})+\gamma_{T^{\prime}}V_{\pi,f}^{g}(\hat{s}_{T^{\prime}})$ using critic $v$ upon reaching state $\hat{s}_{T^{\prime}}$ at the planning horizon $T^{\prime}$. This can repeat multiple times until the planner selects the action sequence $a^{\prime*}_{t:T^{\prime}-1}$ with the highest expected success and passes the first step $a_{t}^{*}$ to actor $\alpha$ which finally outputs the concrete action $a_{t}$. Formally, SimuRA can be seen as solving the following multi-level optimization problem:
@@
$a_{t:T^{\prime}-1}^{\prime*}$ $={\underbrace{\operatorname*{arg\,max}_{a_{t:T-1}^{\prime}}}_{\scriptsize\shortstack{sampled from \\ policy $\tilde{\pi}$}}}\sum_{\hat{s}_{t+1:T^{\prime}}}\ {\underbrace{v(\hat{s}_{T^{\prime}})}_{\text{critic}}}\prod_{k=t}^{T^{\prime}-1}{\underbrace{p_{f}(\hat{s}_{k+1}\mid\hat{s}_{k},a_{k}^{\prime})}_{{\scriptsize\shortstack{world model}}}}$ (Planning) (8)
$a_{t}$ $=\operatorname*{arg\,max}_{a}\ {\underbrace{p_{\alpha}(a\mid\hat{s}_{t},a_{t}^{\prime*})}_{\begin{array}[]{c}\scriptsize\text{actor}\text{}\end{array}}}$ (Acting) (10)
In §4 below, we present an instance of SimuRA with each of these components implemented using pretrained LLMs. While these LLMs alone are often insufficient for many complex agentic tasks, SimuRA’s divide-and-conquer approach combines existing LLM strengths like instruction-following, summarization, reflection, and tool use to allow agentic behavior to emerge. Benefiting from massive web-scale pretraining on next-token prediction $p(x_{t}\mid x_{<t})$, which is formally akin to world modeling, LLMs possess significant potential to serve as world models with natural-language state and simulated action spaces hao2023reasoninglanguagemodelplanning ; hu2023language . We approximately infer the world state $\hat{s}_{t}$ and action $a_{t}$ by sampling from the LLM-based encoder and actor distributions $p_{h}$ and $p_{\alpha}$, respectively. For planning, we optimize over the sampled actions $a_{t:T^{\prime}-1}^{\prime}$ using readily available tree search algorithms like Depth-First Search (DFS) and Monte-Carlo Tree Search (MCTS).
@@ 4 Experiments
Figure 5: LLM-based implementation of SimuRA for web-related tasks (e.g. multi-website QA, flight search, etc). Planner is where we implement our proposed world-model-based planning. We also implement a baseline that simply samples the plan from a language model (i.e., autoregressive planning).
@@
    BrowsingAgent from OpenHands (openhands, ), a representative open-web agent which generates chain-of-thought before selecting an action.
@@
Figure 6: Overview of performance comparison between SimuRA and baselines. The full architecture shows clear advantage over the baseline BrowsingAgent, improving the performance on complex website navigation from 0% to 32.2%. Our proposed world model reasoning for planning also consistently improves over simple planning with autoregressive LLM by up to 124%. *We implement the autoregressive planner with o3-mini and test on the complex website navigation dataset.
@@
Figure 5 presents our implementation when applied to web browsing. We use prompts tailored to the web environments in this example, but plan to extend to other environments and move towards training a single agent model that can act optimally in a wider range of environments and tasks, which is an exciting next step. At each step $t$, the agent receives the observation $o_{t}$ as the HTML-based accessibility tree visible through the browser’s viewport (an example is provided in Appendix A). The agent then uses encoder LLM $h$ to summarizes the observation as $\tilde{s}_{t}\sim p_{h}(\cdot\mid o_{t})$, and then add it to a selective memory of past summaries and simulated actions $\{m(\tilde{s}_{k},a_{k}^{\prime*})\}_{k=1}^{t-1}$ to form the estimated world state $\hat{s}_{t}=[m(\tilde{s}_{1},a_{1}^{\prime*}),\dots,m(\tilde{s}_{t-1},a_{t-1}^{\prime*}),\tilde{s}_{t}]$ for planning. During planning, we sample $M$ simulated actions $a^{\prime}_{t}$ from the policy $\tilde{\pi}$, cluster them into distinct actions, and use the world model $f$ to predict the next summary as $\tilde{s}_{t+1}\sim p_{f}(\cdot\mid\hat{s}_{t},a_{t}^{\prime})$ to form the next state $\hat{s}_{t+1}=[m(\tilde{s}_{1},a_{1}^{\prime*}),\dots,m(\tilde{s}_{t},a_{t}^{\prime}),\tilde{s}_{t+1}]$; this repeats until the planning horizon $T$. To evaluate the terminal state $\hat{s}_{T}$ with critic $v$, we prompt the LLM to generate categorical answers and convert them into numerical scores (e.g., “success” receives a score of 1), and repeat for $N$ times to capture the fine-grained differences between states. Following previous work koh2024tree ; gu2025llmsecretlyworldmodel , we set $M=N=20$ and $T=t+1$, and use DFS as the search algorithm. We implement the planning process using LLM Reasoners hao2024llm , a library for LLM-based complex reasoning using advanced algorithms. After the planner selects the simulated action $a_{t}^{\prime*}$, we update the memory with $m(\tilde{s}_{t},a_{t}^{\prime*})$. For the actor $\alpha$, we additionally include the observation text $o_{t}$ in the prompt to ensure the action grounding. All the prompts are included in Appendix B.
@@
An overview of our results is presented in Figure 6. Across all 3 categories of tasks, our architecture shows a clear advantage over the baseline BrowsingAgent, specifically increase the success rate on complex website navigation from 0% with a popular open-web agent baseline to 32.2%. Our proposed world model reasoning for planning also consistently improves over a matched autoregressive reasoning baseline by up to 124% in task-completion rate. Perhaps surprisingly, autoregressive planning with reasoning LLMs like o3-mini have close-to-zero success rate in Complex Website Navigation, showing that even a strong textual-reasoning model can have limited success with planning for web interactions. All experiments were conducted between November and December 2024 using the best publicly available models and browser environments at that time, except for the o3-mini experiment which was performed in February 2025. While subsequent model updates may yield higher absolute scores on these tasks, our results demonstrate that explicit world-model simulation provides a measurable and repeatable advantage over purely autoregressive planning under identical conditions. In the subsections below, we will introduce the evaluation settings and discuss the results in more detail.
@@ 4.1 Complex Website Navigation
Dataset
Existing benchmarks he2024webvoyagerbuildingendtoendweb have contributed valuable infrastructure for evaluating web agents in diverse real-world settings. However, these datasets are primarily constructed through self-instruct–based task generation followed by human verification, which, while ensuring breadth and realism, offers limited control over task structure and constraint complexity. Therefore, the resulting tasks may follow distributions favored by the language model used for generation, making it difficult to systematically assess how an agent’s reasoning performance scales with task difficulty or compositional complexity. To address this limitation, we introduce FlightQA, a dataset designed for controlled and scalable evaluation of reasoning robustness in live web environments. We focus on flight search, a representative real-world use case that requires multi-step reasoning and precise constraint handling. In this setup, the user requests a flight satisfying a list of constraints (e.g., one-way, from New York to Los Angeles), and the agent must operate an online flight search interface to locate and return a valid option. To systematically evaluate the agent’s reasoning ability, we construct questions with varying number of constraints by iteratively adding to the list, which enables a type of counterfactual analysis that controls for the confounding effect of specific constraint configurations. For instance, an agent capable of robust reasoning should succeed when an additional constraint is added to an existing query, whereas one relying on rote memorization will likely fail.
@@
We ran the experiments and evaluation using gpt-4o between November 24th, 2024 and December 9th, 2024. To evaluate the capabilities of autoregressive reasoning LLMs trained with reinforcement learning, we also included variants of the autoregressive reasoning baseline where the planner is implemented using o1 openai_learning_to_reason and o3-mini openai_o3_mini_2025 and ran them between February 3rd to 5th, 2025. For the environment, we use BrowserGym (workarena2024, ), a popular open-source browser sandbox. We stop each run when the agent provides a response or after the agent takes 30 actions, whichever comes first. We also mark the run as failed when the agent repeats the same action for 3 times consecutively or when the agent causes more than 3 errors while interacting with the browser.
Figure 8: Illustration of the data generation process for the FlightQA dataset. We first prompt a LLM to generate NN lists of $C$ starting constraints (Constraint Generation). Then, we prompt the LLM to iteratively add constraints to the lists one by one, repeating for $K$ times (Constraint Extension). Finally, we prompt the LLM to convert each constraint list into a question in natural language (Question Generation).
@@ Results
We present our Complex Website Navigation results in Table 1. Compared to BrowsingAgent which fails completely in this task, our full architecture improves the accuracy from 0% to 32.2%. Within our architecture, our proposed world-model-based planning shows superior performance over autoregressive reasoning with a 124% improvement (significant at the 0.01 level). The other components in our architecture, which communicate using the structured latent space of model-generated natural language (e.g., observation summary and selective memory), also result in more coherent behavior by reducing the action error rate in BrowsingAgent from 93.3% to 1.1%. However, the autoregressive reasoning still results in frequent repetitions, which is mitigated by world model-based planning (44.4% $\rightarrow$ 18.9%). Perhaps surprisingly, o1 and o3-mini receive close to 0% success rates when serving as autoregressive planners, suggesting that even strong textual-reasoning models can have limited success with planning for web interactions. We do not include BrowsingAgent with o1 and o3-mini as the resulting agent frequently hallucinate answers without interacting with the webpage, which precludes them as viable agents. Those hallucinations, however, fool the LLM evaluator at significant rates, suggesting additional work is needed to secure its robustness.
@@
Table 1: Performance and outcome statistics for the FlightQA dataset. Our architecture increases the accuracy from 0% in OpenHands BrowsingAgent to 32.2%. Reasoning by world model simulation also clearly outperforms autoregressive reasoning by 124%. ** indicates being significantly higher than the second-best method at the statistical significance level of 0.01 ($p<0.01$) based on pairwise t-test. We implement the autoregressive planner with o1 and o3-mini, respectively.
@@
We present our results on General Web Automation in Table 3. Continuing the patterns from previous experiments, our proposed architecture improves over BrowsingAgent by up to 91.7%, while within our architecture, world model reasoning improves over autoregressive reasoning by 21.1%, highlighting the comparative advantage under the given experimental setup. Due to following the environment and evaluator provided by OpenHands, which prioritizes open-web browsing and differs from the standard benchmark setup, the success rates are not directly comparable to those reported in prior work zhou2023webarena . Nevertheless, these results demonstrate the relative advantage of our proposed simulative reasoning architecture.